# Fine-Grained Motion Compression and Selective Temporal Fusion for Neural B-Frame Video Coding

### v1

@@ Abstract
With the remarkable progress in neural P-frame video coding, neural B-frame coding has recently emerged as a critical research direction. However, most existing neural B-frame codecs directly adopt P-frame coding tools without adequately addressing the unique challenges of B-frame compression, leading to suboptimal performance. To bridge this gap, we propose novel enhancements for motion compression and temporal fusion for neural B-frame coding. First, we design a fine-grained motion compression method. This method incorporates an interactive dual-branch motion auto-encoder with per-branch adaptive quantization steps, which enables fine-grained compression of bi-directional motion vectors while accommodating their asymmetric bitrate allocation and reconstruction quality requirements. Furthermore, this method involves an interactive motion entropy model that exploits correlations between bi-directional motion latent representations by interactively leveraging partitioned latent segments as directional priors. Second, we propose a selective temporal fusion method that predicts bi-directional fusion weights to achieve discriminative utilization of bi-directional multi-scale temporal contexts with varying qualities. Additionally, this method introduces a hyperprior-based implicit alignment mechanism for contextual entropy modeling. By treating the hyperprior as a surrogate for the contextual latent representation, this mechanism implicitly mitigates the misalignment in the fused bi-directional temporal priors. Extensive experiments demonstrate that our proposed codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configurations.
@@
B-frame coding exhibits two distinctive characteristics that pose unique challenges. First, it requires estimating bi-directional motion vectors from bi-directional reference frames, leading to higher motion coding costs compared to P-frame coding. Current schemes typically process bi-directional motion vectors by either: (1) concatenating them for joint compression using a single motion auto-encoder [33, 32], or (2) employing a parameter-shared motion auto-encoder for separate compression [34]. However, these coarse-grained motion vector compression methods overlook two fundamental aspects of bi-directional motion characteristics: (a) the temporal correlation asymmetry between forward and backward references leads to distinct bitrate allocation needs and reconstruction quality requirements for bi-directional motion vectors, and (b) the inherent geometric consistency between forward and backward motion vectors can be exploited through joint motion modeling to improve compression efficiency. Second, the quality of predicted bi-directional temporal contexts and temporal priors varies substantially due to the differences in temporal reference correlation and motion vector precision. Existing schemes [33, 32, 34] fused these contexts and priors uniformly through the contextual encoder-decoder and contextual entropy model without quality discrimination, which may propagate prediction errors and ultimately degrade rate-distortion performance. These characteristics suggest that dedicated algorithmic designs for B-frame coding may be necessary to achieve further compression efficiency improvement.
@@ II Related Work
II-A Neural P-Frame Video Coding
The majority of neural video codecs adopt the P-frame coding paradigm [4, 5, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 28, 24, 25, 26, 27, 29, 30, 6, 31, 35, 36, 37, 38, 39, 40, 41]. Lu et al. [4] proposed a pioneering work in this field—DVC [4], which implements key video coding components including motion estimation, motion compression, motion compensation, residual compression, and entropy models with neural networks and optimizes them in an end-to-end manner. Subsequently, Lin et al. [10] proposed to fuse the information of multiple previously compressed motion vectors to reduce the motion coding costs. Agustsson et al. [12] developed a scale-space flow, extending optical flow with a scale parameter to better model uncertainty, along with an enhanced warping operator robust to disocclusions and fast motion. Alternative to optical flow-based approaches, Hu et al. [11] proposed to represent motion with the offsets of deformable convolution and utilize deformable convolution to perform motion compensation. Based on this work, they further proposed to perform deformable convolution-based motion estimation and compensation in a coarse-to-fine manner [5].
@@ II-A Neural P-Frame Video Coding
II-B Neural B-Frame Video Coding
The development of neural B-frame video coding [42, 43, 13, 44, 45, 46, 47, 48, 32, 49, 50, 51, 52, 33, 53, 54] is relatively slower compared with neural P-frame video coding. Based on the need for motion coding, existing neural B-frame coding schemes can be categorized into two types.
Schemes in the first category [42, 43, 13, 44] depend on video interpolation to perform bi-directional temporal prediction without the need for motion coding. Wu [42] et al. designed a contextual video interpolation network that uses bi-directional multi-scale contexts to obtain predicted frames. Following a similar idea, Cheng et al. [13] proposed to adaptively select the frame number in one interpolation loop according to video motion patterns. Xu [44] proposed to select interpolated temporal contexts adaptively using a residual-guided masking encoder. Alexandre et al. [43] introduced a two-layer neural B-frame codec that merges the low-resolution interpolated frames from the base layer and high-resolution interpolated frames from the enhancement layer to generate more accurate predicted frames. Currently, the best interpolation-based neural B-frame video codec [43] has outperformed x265—the industrial software of H.265/HEVC.
Schemes in the second category [45, 46, 47, 48, 32, 49, 50, 51, 52, 33, 53, 54] aim to enhance bi-directional temporal prediction by compressing bi-directional motion information. For example, Yang et al. [45] proposed to perform bi-directional motion vector estimation and compression to obtain bi-directional predicted frames. Subsequently, they employed a residual codec to compress the prediction residuals. Feng et al. [49] proposed to replace the bi-directional optical flows with a voxel flow and replace the bi-directional prediction merging process with the trilinear warping operation. Recently, some schemes [50, 32, 53, 33, 54] proposed to extend conditional coding [27] from P-frame to B-frame. For example, Sheng et al. [33] proposed DCVC-B that predicts bi-directional multi-scale temporal contexts using a bi-directional temporal context mining module. These contexts are then directly fused into a contextual encoder-decoder and a contextual entropy model to reduce temporal redundancy. Although DCVC-B [33] surpassed the reference software of H.265/HEVC standard with random-access configurations, it is difficult to achieve higher B-frame compression performance by simply applying P-frame coding methods to B-frame coding without fully considering the unique characteristics of bi-directional prediction.
@@ III Methodology
The overview of our neural B-frame video codec is illustrated in Fig. 1. It adopts an architecture similar to [33]. To reduce bi-directional motion coding costs and achieve selective utilization of bi-directional temporal predictions, we propose a fine-grained motion compression method and a selective temporal fusion method.
III-A Fine-Grained Motion Compression
@@
III-A1 Interactive Dual-Branch Motion Auto-Encoder
@@
Unlike the scheme that employed a parameter-shared motion auto-encoder for separate bi-directional motion compression [34], our dual-branch motion auto-encoder designs a motion information interaction (MII) module to exploit the inherent geometric consistency between forward and backward motion vectors. As shown in Fig. 2, the motion information interaction module takes intermediate bi-directional motion features (, ) as inputs and outputs (, ). At the start of the MII module, two depth-wise convolution blocks are separately applied to (, ) to generate bi-directional query, key, and value features (), ().
@@
where and are two depth-wise convolution blocks. With our proposed MII module, the information from dual motion branches can be exchanged, thereby exploiting the cross-directional motion dependencies and reducing the motion coding costs.
III-A2 Per-Branch Adaptive Quantization Steps
Adaptive quantization step techniques [35, 36, 38] have proven effective for precise rate-distortion control in neural video coding. While recent work [33] successfully extended this method from P-frame to B-frame coding, it employed identical quantization steps for bi-directional motion vectors. This uniform treatment fails to account for the distinct bitrate allocation needs and reconstruction quality requirements inherent to bi-directional motion vectors. To address this limitation, we propose to integrate independent learnable quantization steps into each branch of both the motion encoder and decoder, as presented in Fig. 2. One quantization step is composed of a global quantization step and a channel-wise quantization step. The global quantization step is a number related to the target bit rate. The global quantization step is a one-dimensional vector, and the length of the vector is the same as the number of channels of the quantized feature.
@@ III-A2 Per-Branch Adaptive Quantization Steps
III-A3 Interactive Motion Entropy Model
@@
Given the quantized bi-directional motion latent representations (, ), we propose an interactive motion entropy model to estimate their probability distributions for arithmetic coding. First, we derive the motion hyperprior and motion temporal prior through:
@@
where and are motion hyper encoder and motion hyper decoder, respectively. is the motion temporal prior extractor [33]. Next, we partition and into four segments via spatial-channel splitting, as illustrated in Fig. 3. To effectively capture bi-directional motion dependencies, when compressing the current motion latent segment, we not only fuse the motion hyperprior, motion temporal prior, and previously compressed latent segments in the same direction, but also fuse the previously compressed latent segments in the other direction.
By interactively fusing bi-directional latent segments as directional priors, our entropy model achieves fine-grained motion dependency modeling and compression efficiency.
III-B Selective Temporal Fusion
@@
Given the reconstructed bi-directional motion vectors, we employ the temporal context mining moulde [31] to extract extract multi-scale temporal contexts from reference features (), , yielding forward contexts and and backward contexts , as illustrated in Fig. 1. These contexts and previously compressed contextual latent representations () are fused into the contextual encoder-decoder and context entropy model to reduce temporal redundancy. Our selective temporal fusion method focuses on how to discriminative utilize these bi-directional temporal contexts and temporal priors.
III-B1 Bi-Directional Weighting-based Context Fusion
@@
Our method first incorporates a bi-directional weighting fusion mechanism to build a discriminative contextual encoder-decoder. As illustrated in Fig. 4, unlike the schemes [33] that blindly fused bi-directional temporal contexts, our discriminative contextual encoder-decoder predicts bi-directional fusion weights to selective fuse the intermediate contextual feature with the bi-directional temporal contexts ().
@@
where is a convolutional layer and refers to element-wise multiplication. Through the end-to-end optimization, the predicted fusion weights adaptively recalibrate the contributions of bi-directional temporal contexts based on their relevance to the target frame, thereby enabling selective temporal fusion that suppresses redundant or noisy temporal contexts.
III-B2 Implicit Alignment-based Prior Fusion
@@
Previous schemes [33] tended to fuse the compressed contextual latent representations (, ) and small resolution of temporal contexts (, ) into the contextual entropy model as a temporal prior . However, these schemes ignore the temporal misalignment of latent representations and the predicted errors of temporal contexts, which fail to optimally exploit temporal correlations. Therefore, our method further incorporates a simple hyperprior-based implicit alignment for the contextual entropy model. Specifically, we first use two depth-wise convolution blocks to generate query, key, and value features () from hyperprior and temporal prior .
@@ IV-A1 Training Datasets
For training, we employ the training split of the Vimeo-90k dataset [57], which comprises 64610 video clips at 448256 resolution. Since each clip contains only 7 consecutive frames, we utilize this dataset for training stages requiring sequence lengths of 3, 5, and 7 frames [33]. For longer sequence requirements (17 and 33 frames), we utilize the extended Vimeo collection introduced in [33], which offers 9000 higher-resolution (640360) sequences with 33 frames per clip. The detailed training strategy can be referred to [33].

### v2

@@ Abstract
With the remarkable progress in neural P-frame video coding, neural B-frame coding has recently emerged as a critical research direction. However, most existing neural B-frame codecs directly adopt P-frame coding tools without adequately addressing the unique challenges of B-frame compression, leading to suboptimal performance. To bridge this gap, we propose novel enhancements for motion compression and temporal fusion for neural B-frame coding. First, we design a fine-grained motion compression method. This method incorporates an interactive dual-branch motion auto-encoder with per-branch adaptive quantization steps, which enables fine-grained compression of bi-directional motion vectors while accommodating their asymmetric bitrate allocation and reconstruction quality requirements. Furthermore, this method involves an interactive motion entropy model that exploits correlations between bi-directional motion latent representations by interactively leveraging partitioned latent segments as directional priors. Second, we propose a selective temporal fusion method that predicts bi-directional fusion weights to achieve discriminative utilization of bi-directional multi-scale temporal contexts with varying qualities. Additionally, this method introduces a hyperprior-based implicit alignment mechanism for contextual entropy modeling. By treating the hyperprior as a surrogate for the contextual latent representation, this mechanism implicitly mitigates the misalignment in the fused bi-directional temporal priors. Extensive experiments demonstrate that our proposed codec achieves an average BD-rate reduction of approximately 10% compared to the state-of-the-art neural B-frame codec, DCVC-B, and delivers comparable or even superior compression performance to the H.266/VVC reference software under random-access configurations.
@@
B-frame coding exhibits two distinctive characteristics that pose unique challenges. First, it requires estimating bi-directional motion vectors from bi-directional reference frames, leading to higher motion coding costs compared to P-frame coding. Current schemes typically process bi-directional motion vectors by either: (1) concatenating them for joint compression using a single motion auto-encoder [57, 7], or (2) employing a parameter-shared motion auto-encoder for separate compression [76]. However, these coarse-grained motion vector compression methods treat bi-directional motions in an undifferentiated manner and thus overlook two fundamental aspects of bi-directional motion characteristics: (a) the temporal correlation asymmetry between forward and backward references leads to distinct bitrate allocation needs and, consequently, different required levels of precision for the reconstruction of bi-directional motion vectors, and (b) the inherent geometric consistency between forward and backward motion vectors can be exploited through joint motion modeling to improve compression efficiency. Second, the quality of predicted bi-directional temporal contexts and temporal priors varies substantially due to the differences in temporal reference correlation and motion vector precision. Existing schemes [57, 7, 76] fused these contexts and priors uniformly through the contextual encoder-decoder and contextual entropy model without quality discrimination, which may propagate prediction errors and ultimately degrade rate-distortion performance. These characteristics suggest that dedicated algorithmic designs for B-frame coding may be necessary to achieve further compression efficiency improvement.
@@ II Related Work
II-A Traditional Video Coding
Traditional video coding has been developed for several decades, leading to the creation of several widely adopted standards, such as H.264/AVC [67], H.265/HEVC [61], and H.266/VVC [4]. These standards follow a similar hybrid coding framework, which typically integrates block-based motion compensated prediction [22, 30, 10], transform [79, 59], quantization [5, 50, 60], entropy coding [62, 50], and in-loop filtering [21, 45]. While each successive generation has delivered substantial improvements in compression performance over its predecessor, the pace of improvement has gradually slowed, indicating a degree of maturity in this classical paradigm. In recent years, the remarkable success of deep learning has motivated the exploration of data-driven approaches, spurring significant interest and rapid progress in neural video coding as a promising complementary direction.
II-B Neural P-Frame Video Coding
The majority of neural video codecs adopt the P-frame coding paradigm [39, 58, 19, 48, 18, 38, 31, 20, 1, 9, 49, 34, 36, 35, 78, 32, 15, 14, 37, 52, 65, 69, 66, 40, 41, 26, 17, 8, 56, 55, 53, 27, 28, 54, 29, 6, 63, 25]. Lu et al. [39] proposed a pioneering work in this field—DVC [39], which implements key video coding components including motion estimation, motion compression, motion compensation, residual compression, and entropy models with neural networks and optimizes them in an end-to-end manner. Subsequently, Lin et al. [31] proposed to fuse the information of multiple previously compressed motion vectors to reduce the motion coding costs. Agustsson et al. [1] developed a scale-space flow, extending optical flow with a scale parameter to better model uncertainty, along with an enhanced warping operator robust to disocclusions and fast motion. Alternative to optical flow-based approaches, Hu et al. [20] proposed to represent motion with the offsets of deformable convolution and utilize deformable convolution to perform motion compensation. Based on this work, they further proposed to perform deformable convolution-based motion estimation and compensation in a coarse-to-fine manner [19].
@@ II-B Neural P-Frame Video Coding
II-C Neural B-Frame Video Coding
The development of neural B-frame video coding [68, 43, 44, 2, 9, 70, 13, 73, 74, 75, 77, 7, 12, 72, 11, 46, 57, 33, 23] is relatively slower compared with neural P-frame video coding. Based on the need for motion coding, existing neural B-frame coding schemes can be categorized into two types.
Schemes in the first category [68, 43, 2, 9, 70] depend on video interpolation to perform bi-directional temporal prediction without the need for motion coding. Wu et al. [68] designed a contextual video interpolation network that uses bi-directional multi-scale contexts to obtain predicted frames. Nguyen et al. [2] adopted a kernel-based, motion-free approach for neural video compression, where video predictions are generated from reference frames via kernel-based interpolation. Cheng et al. [9] proposed to adaptively select the frame number in one interpolation loop according to video motion patterns. Xu [70] proposed to select interpolated temporal contexts adaptively using a residual-guided masking encoder. Alexandre et al. [2] introduced a two-layer neural B-frame codec that merges the low-resolution interpolated frames from the base layer and high-resolution interpolated frames from the enhancement layer to generate more accurate predicted frames. Currently, the best interpolation-based neural B-frame video codec [2] has outperformed x265—an efficient and widely used H.265/HEVC encoding software implementation.
Schemes in the second category [73, 13, 44, 74, 75, 77, 7, 12, 72, 11, 46, 57, 33, 23] aim to enhance bi-directional temporal prediction by compressing bi-directional motion information. For example, Yang et al. [73] proposed to perform bi-directional motion vector estimation and compression to obtain bi-directional predicted frames. Subsequently, they employed a residual codec to compress the prediction residuals. Feng et al. [12] proposed to replace the bi-directional optical flows with a voxel flow and replace the bi-directional prediction merging process with the trilinear warping operation. Recently, some schemes [72, 7, 13, 33, 57, 23] proposed to extend conditional coding [26] from P-frame to B-frame. For instance, Gao et al. [13] introduced MaskCRT-B, a B-frame coding method based on a masked conditional residual transformer. It employs conditional residual coding and a bi-directional adaptive fusion module to enhance the prediction of occluded and disoccluded regions. Building upon this work, NguyenQuang et al. [44] developed a fast online motion resolution adaptation method. This method employs lightweight classifiers to determine the optimal motion downsampling factor, thereby addressing failures in large motion prediction. Recently, Sheng et al. [57] proposed DCVC-B that predicts bi-directional multi-scale temporal contexts using a bi-directional temporal context mining module. These contexts are then directly fused into a contextual encoder-decoder and a contextual entropy model to reduce temporal redundancy. Although DCVC-B [57] surpassed the reference software of H.265/HEVC standard with random-access configurations, it is difficult to achieve higher B-frame compression performance by simply applying P-frame coding methods to B-frame coding without fully considering the unique characteristics of bi-directional prediction.
@@ III Methodology
We begin with a framework overview, highlighting the relationship to the baseline DCVC-B [57], and then detail our proposed fine-grained motion compression and selective temporal fusion methods.
III-A Framework Overview
The architecture of our neural B-frame video codec is illustrated in Fig. 1. While following a similar architecture to DCVC-B [57], our work introduces key innovations to address specific limitations in B-frame coding. The framework retains several components from the baseline, including bi-directional motion estimation, temporal context mining, and frame generation. Our main contributions focus on substantially improving the motion compression and contextual compression. First, to reduce bi-directional motion coding costs, we propose a fine-grained motion compression method that replaces the original coarse-grained motion compression with an interactive dual-branch auto-encoder, per-branch adaptive quantization, and an interactive motion entropy model. Second, to enable discriminative utilization of temporal contexts and priors in the contextual compression, we develop a selective temporal fusion method that supersedes the original uniform fusion strategy through bi-directional weighting and implicit prior alignment. These key innovations are designed to tackle the unique challenges of B-frame coding, leading to substantial gains in compression efficiency.
III-B Fine-Grained Motion Compression
@@
III-B1 Interactive Dual-Branch Motion Auto-Encoder
@@
Unlike the scheme that employed a parameter-shared motion auto-encoder for separate bi-directional motion compression [76], our dual-branch motion auto-encoder designs a motion information interaction (MII) module to exploit the inherent geometric consistency between forward and backward motion vectors. As shown in Fig. 2, the motion information interaction module takes intermediate bi-directional motion features (, ) as inputs and outputs (, ). At the start of the MII module, two depth-wise convolution blocks are separately applied to (, ) to generate bi-directional query (Q), key (K), and value (V) features (), ().
@@
where and are two depth-wise convolution blocks. refers to channel-wise concatenation. With our proposed MII module, the information from dual motion branches can be exchanged, thereby exploiting the cross-directional motion dependencies and reducing the motion coding costs.
III-B2 Per-Branch Adaptive Quantization Steps
Adaptive quantization step techniques [27, 28, 29] have proven effective for precise rate-distortion control in neural video coding. While recent work [57] successfully extended this method from P-frame to B-frame coding, it employed identical quantization steps for bi-directional motion vectors. This uniform treatment fails to account for the distinct bitrate allocation needs and reconstruction quality requirements inherent to bi-directional motion vectors. To address this limitation, we propose to integrate independent learnable quantization steps into each branch of both the motion encoder and decoder, as presented in Fig. 2. One quantization step is composed of a global quantization step and a channel-wise quantization step. The global quantization step is a number related to the target bit rate. The channel-wise quantization step is a one-dimensional vector, and the length of the vector is the same as the number of channels of the quantized feature.
@@ III-B2 Per-Branch Adaptive Quantization Steps
III-B3 Interactive Motion Entropy Model
@@
Given the quantized bi-directional motion latent representations (, ), we propose an interactive motion entropy model to estimate their probability distributions for arithmetic coding. First, we derive a motion hyperprior through a motion hyper encoder and a motion hyper decoder :
Then, we use a motion temporal prior extractor [57] to leverage the temporal correlation in motion vectors.
The extractor , composed of four convolutional layers with downsampling, takes the concatenated bi-directional motion vectors (, ) as inputs and generates a motion temporal prior . The output has the same spatial resolution and channel number as the motion latent representations (, ), allowing it to serve as a direct motion temporal prior in the motion entropy model.
@@
Next, we partition and into four segments via spatial-channel splitting [27, 28, 29, 57, 56, 54], as illustrated in Fig. 3. To effectively capture bi-directional motion dependencies, when compressing the current motion latent segment (), we not only fuse the motion hyperprior , motion temporal prior , and previously compressed latent segments () in the same direction, but also fuse the previously compressed latent segments in the other direction ().
where denotes a Laplacian distribution. Our entropy model achieves fine-grained motion dependency modeling by interactively fusing bi-directional latent segments as priors. This strategy, while introducing a partial sequential dependency that limits full parallel decoding across the two directions of motion streams, effectively exploits the inherent geometric correlations in bi-directional motion vectors. By leveraging cross-directional motion prior information, it achieves more accurate motion entropy estimation and enhanced compression performance.
III-C Selective Temporal Fusion
@@
Given the reconstructed bi-directional motion vectors, we employ the temporal context mining moulde [53] to extract multi-scale temporal contexts from reference features (), yielding forward contexts and and backward contexts , as illustrated in Fig. 1. These contexts and previously compressed contextual latent representations () are fused into the contextual encoder-decoder and context entropy model to reduce temporal redundancy. Our selective temporal fusion method focuses on how to discriminatively utilize these bi-directional temporal contexts and temporal priors.
III-C1 Bi-Directional Weighting-based Context Fusion
@@
Our method first incorporates a bi-directional weighting fusion mechanism to build a discriminative contextual encoder-decoder. As illustrated in Fig. 4, unlike the schemes [57] that blindly fused bi-directional temporal contexts, our discriminative contextual encoder-decoder predicts bi-directional fusion weights to selectively fuse the intermediate contextual feature with the bi-directional temporal contexts ().
@@
where is a convolutional layer. Through the end-to-end optimization, the predicted fusion weights adaptively recalibrate the contributions of bi-directional temporal contexts based on their relevance to the target frame, thereby enabling selective temporal fusion that suppresses redundant or noisy temporal contexts.
III-C2 Implicit Alignment-based Prior Fusion
@@
Previous schemes [57] tended to fuse the compressed contextual latent representations (, ) and small resolution of temporal contexts (, ) into the contextual entropy model as a temporal prior . However, these schemes ignore the spatial misalignment of latent representations and the predicted errors of temporal contexts, which fail to optimally exploit temporal correlations. Therefore, our method further incorporates a simple hyperprior-based implicit alignment for the contextual entropy model. Specifically, we first use two depth-wise convolution blocks to generate query, key, and value features () from hyperprior and temporal prior .
@@ IV-A1 Training Datasets
Following previous works [26, 27, 28, 29, 57, 44, 17, 7, 56, 54], for training, we employ the training split of the Vimeo-90k dataset [71], which comprises 64610 video clips at 448256 resolution. Since each clip contains only 7 consecutive frames, we utilize this dataset for training stages requiring sequence lengths of 3, 5, and 7 frames [57]. For longer sequence requirements (17 and 33 frames), we utilize the extended Vimeo collection introduced in [57], which offers 9000 higher-resolution (640360) sequences with 33 frames per clip. The detailed training strategy can be referred to [57].